OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)? (x - 5)2 + (y - 7)2 = 37 (x - 5)2 + (y - 7)2 = 74 (x + 5)2 + (y + 7)2 = 74 (x + 5)2 + (y + 7)2 = 37

geerky42 (geerky42):

Form of circle equation is \[(x-h)^2+(y-k)^2=r^2\]

geerky42 (geerky42):

Where \((h,k)\) is center and \(r\) is radius

geerky42 (geerky42):

You can use distance formula (without square root) to find \(r^2\): \[d^2 = (x_1-x_2)^2+(y_1-y_2)^2\]

geerky42 (geerky42):

You can have \((x_1,y_1) = (-5,-7)\) and \((x_2,y_2) = (0,0)\)

geerky42 (geerky42):

So you would have this after plugging in: \[d^2 = (-5-0)^2+(-7-0)^2 = \cdots~?\]